OpenStudy (anonymous):

How do I determine the ordered pair for the y-intercept of the line 3y-2x=-15

OpenStudy (anonymous):

its actually pairs

OpenStudy (anonymous):

Get it into the form y = mx + b where m is the slope and b is the y-intercept. 3y - 2x = -15 3y = 2x - 15 y = 2/3x - 5 The y-intercept now is the point at which the line or graph intercepts the y-axis. That means that the x in the ordered pair will always be 0. Then, Plug in the y-intercept for the y value. (0, -5)

OpenStudy (anonymous):

so i can use the slope to get the next ordered pair?